OpenStudy (anonymous):

The average speed of one airplane is 200 mph faster than that of an automobile. If the automobile starts at 3 a.m. and the airplane starts from the same place at 11 a.m., they will have traveled the same number of miles by 1:00 p.m. Find the rate of the automobile.

OpenStudy (anonymous):

r t = d Let x be the speed of the automobile. From the problem statement: 8 x + 2 x = 2 (200 + x) x = 50 mph